If tan a= 2k - x÷k√3 and tan b= x√3 ÷ 2k -x then prove that (a-b)=30°.

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
Use 
tan(a-b) = tan A - tanB/(1+tanA tanB)
Put tanA and tanB
 
You will get
tan(a-b) = 1/sqrt(3)
Hence
a - b = 30°
 
Regards
Arun (askIITians forum expert)